OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (adamaero):

put g(x) into f(x) wherever x is

OpenStudy (anonymous):

f(g(x)= f(10x-6)= -4(10x-6) +7. Easy :)

OpenStudy (adamaero):

x in function f(x) becomes function g(x)

OpenStudy (anonymous):

f(g(x)) = -4(10x-6) + 7 = -40x+24 + 7 = -40x + 31

OpenStudy (adamaero):

shouldn't have just given the answer

OpenStudy (adamaero):

@sapphire2018 @kevin4762

OpenStudy (anonymous):

most people come here to find answers for their homework I don't think they actually want to really learn anything. otherwise they wouldn't have asked on here.

OpenStudy (anonymous):

ah, you are right, silly me :p, adamaero

OpenStudy (anonymous):

one should screw the answer up :)

OpenStudy (anonymous):

mixing other equations in hehehe :)

OpenStudy (anonymous):

next time I put a link to a graph instead lol

OpenStudy (adamaero):

lolololol @sapphire2018